# code[Code][List of Codes]

### v1

@@ V-B Main Results and Analysis
TouchGuide can explicitly steers base policy in the action space using touch guidance, and it achieves clear performance gains over current SOTA visuo-tactile policies (Q2). Existing methods primarily focus on reacting promptly to changes in tactile signals (e.g., RDP [65] and PolicyConsensus [9]). However, in many tasks, the initial contact often plays a decisive role in task success. For tasks such as lock opening and shoe lacing, small errors at first contact can cause the key to drop or become severely misaligned, driving the system into states that are entirely unfamiliar to the policy, which is particularly detrimental for imitation learning. In contrast, our method leverages a Contact Physical Model to assess the feasibility of the noisy actions proposed by the base policy and uses the resulting feasibility scores to steer the policy at inference time. This paradigm of evaluating the physical consistency between noisy actions and tactile signals is crucial for producing appropriate contacts. Table I supports this claim, showing that TouchGuide (36.2%) achieves a noticeable performance improvement over other SOTA visuo-tactile policies (e.g., RDP [65] (30.3%) and PolicyConsensus [9] (24.7%)).
TouchGuide demonstrates strong generalization, and can be applied across different robots, tactile modalities, and policies (Q3). Unlike prior visuo-tactile policies, TouchGuide augments a visuomotor policy with touch guidance in a fully decoupled manner, requiring no retraining to the base policy. As a result, it can inherit the strong capability of the base policy. For example, TouchGuide built on $\pi_{0.5}$ achieves solid performance on a high-precision task such as lock opening with only 20 demonstrations. Moreover, TouchGuide achieves strong performance across different robots (e.g., Bi-ARX5 and Flexiv Rizon4), different policies (e.g., DP and $\pi_{0.5}$), and different tactile modalities (e.g., force and tactile image). Detailed results are reported in Table I. More baseline failure cases are provided in Appendix -C.
@@ V-D System Comparison and User Study
To evaluate the usability of our TacUMI, we recruited five volunteers working in robotics (two had prior experience collecting demonstrations using a handheld gripper, while three had no such experience). Each volunteer collected 20 valid demonstrations using VR Teleop and UMI with three different tracking methods (e.g., SLAM-based, AR motion tracker, and TacUMI). We used the following metrics: Attempts (the number of trials required to obtain the target number of demonstrations), Length (the dataset duration excluding preparation and post-processing time, which helps assess system usability), Valid Rate (Valid Data / Attempts), and Satisfaction (each volunteer’s subjective usability rating of the system). As shown in Table V, although Teleop achieves a high valid rate, the tactile feedback is not directly available, which leads to excessively long demonstrations and may degrade policy performance. SLAM-based tracking yields a very low valid rate. The AR Motion Tracker requires wearing an AR headset, making it unsuitable for prolonged data collection, and occlusions can easily cause tracking loss if the user is not constantly attentive during collection. TacUMI demonstrates excellent usability and receives high satisfaction ratings from the volunteers. For a comparative analysis with alternative data collection systems, please refer to Appendix -D1.

### v2

@@ V-B Main Results and Analysis
TouchGuide explicitly steers the base policy in action space via touch guidance, yielding performance gains over current SOTA visuo-tactile policies (Q2). Existing methods mainly target rapid reactions to tactile changes (e.g., RDP [65] and PolicyConsensus [9]); however, RDP relies on low-dim tactile embeddings that discard in-hand state cues, hindering fine-grained manipulation. In many tasks, initial contact is decisive: for Lock Opening and Shoe Lacing, small errors at first contact can drop the key or severely misalign it, pushing the system into out-of-distribution states that are especially problematic for imitation learning. In contrast, our method uses a CPM to assess the feasibility of noisy actions proposed by the base policy and steers the policy at inference time using the resulting feasibility scores. Evaluating physical consistency between actions and tactile signals is crucial for producing appropriate contacts. Table I supports this claim: TouchGuide (36.2%) improves over other SOTA visuo-tactile policies, including RDP [65] (30.3%) and PolicyConsensus [9] (24.7%). Meanwhile, TouchGuide adds negligible inference time overhead (see Appendix -F3).
TouchGuide demonstrates strong generalization, and can be applied across different robots, tactile modalities, and policies (Q3). Unlike prior visuo-tactile policies, TouchGuide augments a visuomotor policy with touch guidance in a fully decoupled manner, requiring no retraining to the base policy. As a result, it can inherit the strong capability of the base policy. For example, TouchGuide built on $\pi_{0.5}$ achieves solid performance on a high-precision task such as Lock Opening with only 20 demonstrations. Moreover, TouchGuide achieves strong performance across different robots (e.g., Bi-ARX5 and Flexiv Rizon4), different policies (e.g., DP and $\pi_{0.5}$), and different tactile modalities (e.g., force and tactile image). Detailed results are reported in Table I. More baseline failure cases are provided in Appendix -D.
@@ V-D System Comparison and User Study
To evaluate the usability of our TacUMI, we recruited five volunteers working in robotics (two had prior experience collecting demonstrations using a handheld gripper, while three had no such experience). Each volunteer collected 20 valid demonstrations using VR Teleop and UMI with three different tracking methods (e.g., SLAM-based, AR motion tracker, and TacUMI) on Lock Opening task. We used the following metrics: Attempts (the number of trials required to obtain the target number of demonstrations), Length (the dataset duration excluding preparation and post-processing time, which helps assess system usability), Valid Rate (Valid Data / Attempts), and Satisfaction (each volunteer’s subjective usability rating of the system). As shown in Table V, although Teleop achieves a high valid rate, the tactile feedback is not directly available, which leads to excessively long demonstrations and may degrade policy performance. SLAM-based tracking yields a very low valid rate. The AR Motion Tracker requires wearing an AR headset, making it unsuitable for prolonged data collection, and occlusions can easily cause tracking loss if the user is not constantly attentive during collection. TacUMI demonstrates excellent usability and receives high satisfaction ratings from the volunteers. More comparisons with existing data collection systems are provided in Appendix -E1.